# Cosmological Zoom-In Simulation of Odd Radio Circles as Merger-Driven Shocks in Galaxy Groups

### v1

@@ 3.2 X-ray properties of the circumgalactic medium
First we will focus on the evolution of the hot gas component and and its inferred X-ray properties. Fig. 2 demonstrates them for two different times – shortly before and after the merger. The first column displays the same X-ray maps as in the first two rows of Fig. 1 in the orthogonal projection, while the second and third column show the radial profiles of the mass-weighted median of the electron density and temperature for the gas of both merging halos, respectively. Note that as a filter for the X-ray emitting gas we applied density and temperature cuts of and , respectively. After the first pericenter passage, the gaseous component of the galaxies is effectively shredded, while the dark matter component still maintains two significant overdensities. Therefore we center the profiles on the two potential minima inferred from the underlying dark matter distribution, which mixes on longer timescales. To minimize contamination by the other halo, the radial profiles are only computed in 3D cones with an opening angle of , pointing away from each other. These cones are indicated by the red and blue shaded regions in the X-ray map on the left in Fig. 2, where the respective color is also adopted for both galaxies in the adjacent radial profiles.
@@
Before the merger, at , the density and temperature profiles of both galaxies suggest two relaxed galaxies, where the gas is in hydrostatic equilibrium. The central density, and hence total mass of one halo (blue line) is approximately two times larger than the other (the exact merger mass ratio is ). At – after the initial pericenter passage – the gas properties exhibit considerable differences both in density and temperature. While the initially less massive galaxy (red line) develops a pronounced core, the other displays a significant drop in density at its inner few kpc, corresponding to a peak in temperature for this galaxy. Both halos exhibit a large temperature jump around a radial distance of , after which the temperature in the CGM of both galaxies drops again. This is coinciding with the position of the shock at this moment, also evident by the drop of the X-ray flux visible in the map on the left of Fig. 2. The pericenter passage led to massive stripping of gas from the central regions of both galaxies, visible by the strong decrease in density compared to the picture before. However, one galaxy (red line) still maintains a much more dense core compared to the other one, whose central density is almost an order of magnitude lower.
@@
This has interesting consequences for the X-ray brightness of both galaxies. The gas luminosity due to thermal X-ray emission is only weakly dependent on temperature, while being quite sensitive to the gas density. Since one galaxy here (red line) has a significantly larger central density, it is consequently outshining the second galaxy. X-ray observations of ORC Physalis revealed that the X-ray flux of the system is centered on one of its two merging galaxies, while the other does not display any significant central emission (c.f. Fig. 2 by Koribalski et al. 2024b). Our simulation manages to reproduce this behavior, considering that the densities of the galaxies are overall very low after the merger, which could be the reason why one halo is still detected, while the other is simply too dilute to be visible in X-ray. Forthcoming deeper X-ray data accumulated for Physalis by (Khabibullin et al., in prep) should allow one to test this prediction.
@@ 3.3 Mock radio observations
In Fig. 3 we present mock images of the radio flux density due to synchrotron emission at four different frequencies and the resolved spectral indices of the simulated ORC. The orientation of the sightline is orthogonal to the merger axis and we choose the time of best morphological resemblance to observed counterparts ( after the merger event). We shift the object from the simulation redshift of to , in order to be consistent with the distance to the observed ORC Physalis (Koribalski et al. 2024b)222Strictly speaking, this is slightly affecting the inferred radio properties of the system, since we include redshift dependent energy losses of CRs via inverse Compton scattering with CMB photons. The spectra are therefore biased towards higher spectral indices here, but the changes to the results are insignificant.. The frequencies and gaussian smoothing kernel widths are chosen such that they coincide with those of currently operating radio telescopes: (, LOFAR), (, GMRT), (, ASKAP) and (, ASKAP). The respective telescope and observer frequency is written in each panel, while the scale in the upper left panel is indicating a physical size of . Similarly to other current works, our simulation is underestimating the local magnetic field strength since we lack the ultra-high resolution necessary to resolve the turbulent dynamo and hence the main driver for efficient magnetic field amplification (Steinwandel et al. 2020), as well as astrophysical seeding mechanisms (e.g. Beck et al. 2013; Garaldi et al. 2021). In order to still provide realistic mock images of the synchrotron brightness – which in turn is heavily relying on the magnetic field strength – we calculate the synchrotron emissivity by assuming equipartition between turbulent and magnetic pressure. This assumption implies a magnetic field strength of , where and are the local gas density and its turbulent velocity component, respectively (Zhou et al. 2024; Böss et al. 2024, c.f. Sect. 4 for details). Recent simulations indicate that this saturation of the turbulent dynamo is valid in regions with strong shocks, where the generation and amplification of magnetic fields is particularly efficient (Steinwandel et al. 2024; Zhou et al. 2024).
@@
The diffuse structure visible at low frequencies is successively reduced at increasing frequencies. This is due to the strong energy dependence of the cooling time of cosmic rays. Therefore we see high energy cosmic ray electrons predominantly only at shocks, where they are currently accelerated. Electrons at lower energies, however, will maintain their energy longer after being shock-accelerated and hence will continue to emit in radio downstream of the shock. This is the reason for the difference between the two the spectral index maps and in the right column of Fig. 3, which indicate the spectral slope of the radio flux density at low and high frequencies, respectively (). In these images we applied a radial mask, in order to reduce the noise in the outskirts (indicated by the dashed circle in the upper left panel). The spectral slope at low frequencies is generally flatter (with maximum values of about at the shock front), than at high frequencies (about at shock), where the bulk of cosmic ray electrons has already cooled off to lower energies, causing a steepening of the spectrum. This aspect, as well as a comparison to observed counterparts will be described in Sect. 3.4.
@@ 3.4 Emission spectrum and total radio power
According to the formation model we are analyzing in this paper, the radio feature that we identify as an ORC is a merger shock around a massive galaxy or galaxy group, that is triggered along the axis of a major merger. In this scenario, the total radio luminosity would be the highest shortly after the merger event, since that would be the time when the shock can inject many CR electron populations efficiently into the relatively dense medium at small radii. If we assume that CRs get accelerated from the thermal pool according to DSA, the powerlaw index of the injected CR’s energy spectrum will depend only on the Mach number . Rewriting in terms of the spectral index of the emission spectrum yields:
@@
In the limit of a strong shock (), the spectral index will be , while weaker shocks will generate steeper spectra (). After the initial central merger, the shock will travel radially outwards with decreasing effective Mach numbers, producing increasingly steep spectra at the front.
@@
In Fig. 4 we present the time evolution of the ORC’s total luminosity produced by synchrotron emission. For this purpose we extracted the particles around the expanding shock front at each time step and calculated their cumulative radio brightness at 50 frequencies equidistant in space according to Eq. A1 by Böss et al. (2023a). Both panels show the total brightness on the y-axis, while the x-axis represents time and frequency in the left and right panel, respectively. On the left, different colors identify varying frequencies according to the colorbar, while the solid black line is highlighting specifically , namely the frequency in which the ORC Physalis (Koribalski et al. 2024b) was observed. The vertical dashed lines indicate the three times at which the simulation was displayed in the rows of Fig. 1 (). Shortly after , the merger triggers the shock – evident by the sudden jump in luminosity – and reaches its maximum by around . Afterwards the total power is steadily decreasing in all frequencies, with the decline being stronger for higher frequencies. This effect becomes even more apparent on the right side of Fig. 4. This panel shows the cumulative emission spectrum of the simulated ORC at four successive times, which are approximately apart, starting from the time of maximum radio brightness at (yellow line). This early-on, we see a relatively flat spectral index of in the lower frequency range. Moving on to later times, the spectrum is rapidly decreasing in total luminosity, with further spectral steepening at higher frequencies. This is due to two effects: while the shock is traveling outwards its effective Mach number is decreasing, which results in a lower acceleration efficiency in the DSA framework implemented in our code. Newly injected CRs therefore display increasingly steep spectra with time (c.f. Eq. 1), while simultaneously the norm of the spectrum is systematically decreasing. The latter feature is also influenced by the decline in local gas density and magnetic field strength with radius, since these circumstances lead to less gas with new CR populations and lower synchrotron emissivity, respectively. For a fixed time step, the steeper slope at high frequencies is caused by radiation losses, since the energy loss time scale of electrons due to Synchrotron emission and inverse Compton scattering is strongly dependent on their energy. This leads to a more rapid energy loss of CR electrons at higher energy bins, which is reflected in the steeper emission spectrum at higher frequencies, since the contribution to the spectral emission of a charged particle in a magnetic field is closely peaked around an energy dependent maximum (Ginzburg 1979).
@@
Taking the spectral index at the earliest time () as a proxy for the overall maximum over time, we can easily estimate the respective Mach number by rearranging Eq. 1 for . This yields that the maximum Mach number of the shock in our simulation is , which is comparable to similar merger events in galaxy clusters. The maximum power in radio at is , which is about three orders of magnitude lower that what is measured at the ORC Physalis (Koribalski et al. 2024b).
3.5 Polarization
@@
Synchrotron radiation of electrons will be polarized with a polarization angle , i.e. the angle between the electric field vector of the associated radiation, and a chosen unit vector in the sky plane, which is defined to be orthogonal to the line of sight. Since the polarization angle is depending only on local properties, the so called polarization fraction will be relatively high in case of a singular, relatively small emitting cosmic ray cloud where the local magnetic field is approximately uniform. But as the number of contributing electrons along the line of sight is increasing, the fractional polarization of the observed radiation will consequently decrease because the direction of the magnetic field is not homogeneous. Therefore, a polarization analysis of radio observations can be a powerful tool in order to dissect the physical processes relevant in ORCs. Since our simulation is self-consistently evolving cosmic ray electrons, we can provide polarization information of the simulated ORC and compare it with observed counterparts for the first time.
@@
The polarization state of electromagnetic radiation at a given frequency is commonly described by the four Stokes parameters and , because they can be inferred directly from observed quantities. Generally, synchrotron radiation is elliptically polarized. But in the limit of ultra-relativistic electrons, the light cone of radiation – which is defined as the solid angle around the velocity vector in which an observer is receiving a significant signal – becomes sufficiently narrow. In this case, the observer will receive only ?flashes? of synchrotron radiation when the electron is moving almost exactly towards the observer. The associated electric field vector is hence only drawing a line in the sky plane (instead of an ellipse), which is equivalent to linear polarization (Ginzburg 1979; Longair 2011). In this limit, is vanishing and only three Stokes parameters remain. Assuming the line of sight is along the -direction, is defined as the total intensity received by the observer integrated along the line of sight:
@@
where is the synchrotron emissivity at a given frequency and position . Meanwhile and , which are actually carrying the polarization information, are defined by
@@
where is the polarized synchrotron emissivity. Note that the polarization angle will be subject to Faraday rotation along the line of sight. The polarization angle therefore needs to be adjusted at each contributing particle along the line of sight according to the traveled distance by
@@
where is the observed wavelength, RM the rotation measure, the local electron number density and the component of the magnetic field vector along the line of sight. The polarization angle is depending only on the magnetic field components in the sky plane and is given by
Finally, the polarization fraction can be obtained via
@@
Generally, two factors will have an influence on the degree of polarization of synchrotron radiation: i) the energy spectrum of the underlying cosmic ray electron population and ii) the turbulence of the magnetic field and hence of the carrying intergalactic medium. In case of a uniform (i.e. non-turbulent) magnetic field, the energy-integrated polarization fraction only depends on the powerlaw index of the CR electron energy distribution, which can be rewritten in terms of the spectral index of the emission spectrum (c.f. Eq. 5.46 by Ginzburg 1979):
@@
The subscript 0 is indicating here that the Stokes parameters from Eq. 7 were integrated over the whole energy range in this case. Although we are regarding only spectral values in this Section, it is still instructive to keep this limit in mind, since it is providing us with a sense of what to expect for a given energy distribution. If the value for the fractional polarization measured in the simulation is significantly lower than that, it will be likely due to turbulence of the magnetic field along the line of sight, since this will result in an increasingly diverse distribution of polarization angles, which can further ?fan out? due to Faraday rotation.
@@
In Fig. 5 we present the polarization fraction of the simulated ORC at a frequency of given by Eq. 7. Each quantity ( and ) was smoothed according to the resolution of ASKAP at (, c.f. Sect. 3.3), where we assume the same distance as to the ORC Physalis at (Koribalski et al. 2024b). The time, scale, orientation and circular mask of the map is the same as in Fig. 3. For the most part, the simulated ORC has a degree of polarization between . We have analysed the rotation measure for the simulation and found that it is negligible outside the halo center, which means that polarization is dominated by the emission at the shock ring (c.f. Fig. 8). Judging by their emission spectrum, CRs are injected quite efficiently in our simulation and yield a spectral index close to the DSA limit of (c.f. Sect. 3.4). The fact that we arrive at a much lower mean polarization degree than the expectancy value given by Eq. 8, is an indicator for the turbulence that is present inside the shocked plasma along the line of sight, as explained above. The sporadic white spots in Fig. 5 of deviantly high degree of polarization in comparison, are sightlines with singular cosmic ray populations, which are therefore not contaminated with additional emissions along the line of sight. This is mostly a numerical artifact, which is also evident by the overall ?popcorn-like? appearance of the map – although we are already working with very high resolution here, the shock finder (which is highly resolution dependent) is occasionally failing to pinpoint the location of the shock smoothly (c.f. Sect. 3.6). Additionally, the injection of CRs is strongly dependent on the evaluated Mach number, as well as the estimated shock obliquity, that is determining the local acceleration efficiency, which can further add uncertainties to the simulated polarization. We point out, however, that this has no significant impact on the the overall value of the polarization fraction across the shock ring. Fig. 5 therefore qualifies for a comparison with observed ORCs, which will be carried out in the Discussion (Sect. 4).

### v2

@@ 3.2 X-ray properties of the circumgalactic medium
First, we focus on the evolution of the hot gas component and its inferred X-ray properties. Fig. 2 demonstrates them for two different times – shortly before and after the merger. The first column displays the same X-ray maps as in the first two rows of Fig. 1 in the orthogonal projection, while the second and third column show the radial profiles of the mass-weighted median of the electron density and temperature for the gas of both merging haloes, respectively. Note that as a filter for the X-ray emitting gas, we applied density and temperature cuts of and , respectively. After the first pericentre passage, the gaseous component of the galaxies was effectively shredded, while the dark matter component maintains two significant overdensities. Therefore, we centred the profiles on the two potential minima inferred from the underlying dark matter distribution, which mixes on longer timescales. To minimise contamination by the other halo, the radial profiles were only computed in 3D cones with an opening angle of , pointing away from each other. These cones are indicated by the red- and blue-shaded regions in the X-ray map on the left in Fig. 2, where the respective colour is also adopted for both galaxies in the adjacent radial profiles.
@@
Before the merger, at , the density and temperature profiles of both galaxies suggest two relaxed galaxies, where the gas is in hydrostatic equilibrium. The central density, and hence total mass of one halo (blue line), is approximately two times higher than the other (the exact merger mass ratio is ). At – after the initial pericentre passage – the gas properties exhibit considerable differences in both density and temperature. While the initially less massive galaxy (red line) develops a pronounced core, the other displays a significant drop in density at its inner few kiloparsecs, corresponding to a peak in temperature for this galaxy. Both haloes exhibit a large temperature jump around a radial distance of , after which the temperature in the circumgalactic medium (CGM) of both galaxies drops again. This coincides with the position of the shock at this moment, also evident by the drop of the X-ray flux visible in the map on the left of Fig. 2. The pericentre passage led to massive stripping of gas from the central regions of both galaxies, visible by the strong decrease in density compared to the picture before. However, one galaxy (red line) maintains a much denser core compared to the other one, whose central density is almost an order of magnitude lower.
@@
This has interesting consequences for the X-ray brightness of both galaxies. The gas luminosity due to thermal X-ray emission is only weakly dependent on temperature while being quite sensitive to gas density. Since one galaxy here (red line) has a significantly higher central density, it consequently outshines the second galaxy. X-ray observations of ORC Physalis revealed that the X-ray flux of the system is centred on one of its two merging galaxies, while the other does not display any significant central emission (cf. Fig. 2 by Koribalski et al. 2024b). Our simulation managed to reproduce this behaviour, considering that the galaxy densities overall remain very low after the merger. This could explain why one halo is still detected, while the other is too dilute to be visible in X-ray. Forthcoming deeper X-ray data accumulated for Physalis by XMM-Newton (Khabibullin et al., in prep) should allow one to test this prediction.
@@ 3.3 Mock radio observations
In Fig. 3 we present mock images of the radio flux density due to synchrotron emission at four different frequencies and the resolved spectral indices of the simulated ORC. The orientation of the sight line is orthogonal to the merger axis, and we chose the time of best morphological resemblance to observed counterparts ( after the merger event). We shifted the object from the simulation redshift of to , in order to be consistent with the distance to the observed ORC Physalis (Koribalski et al. 2024b)444Strictly speaking, this slightly affects the inferred radio properties of the system, since we include redshift-dependent energy losses of CRs via inverse Compton scattering with photons of the cosmic microwave background. The spectra are therefore biased towards higher spectral indices here. Nevertheless, the qualitative changes to the results are insignificant.. The frequencies and Gaussian smoothing kernel widths were chosen such that they coincide with those of currently operating radio telescopes: (, LOFAR), (, GMRT), (, ASKAP) and (, ASKAP). The respective telescope and observer frequency is written in each panel, while the scale in the upper left panel indicates a physical size of . Similarly to other current studies, our simulation underestimates the local magnetic field strength since we lack the ultra-high resolution necessary to resolve the turbulent dynamo and hence the main driver for efficient magnetic field amplification (Steinwandel et al. 2020), as well as astrophysical seeding mechanisms (e.g. Beck et al. 2013; Garaldi et al. 2021). In order to provide realistic mock images of the synchrotron brightness – which in turn heavily relies on the magnetic field strength – we calculated the synchrotron emissivity by assuming equipartition between turbulent and magnetic pressure throughout this work. This assumption implies a magnetic field strength of , where and are the local gas density and its turbulent velocity component, respectively (Zhou et al. 2024; Böss et al. 2024, see Sect. 4 for details). Recent simulations indicate that this saturation of the turbulent dynamo is valid in regions with strong shocks, where the generation and amplification of magnetic fields is particularly efficient (Steinwandel et al. 2024; Zhou et al. 2024).
@@
The diffuse structure visible at low frequencies is successively reduced at increasing frequencies. This is due to the strong energy dependence of the CR cooling time. Therefore, we see high-energy CR electrons predominantly only at shocks, where they are currently accelerated. Electrons at lower energies, however, maintain their energy longer after being shock-accelerated and hence continue to emit in radio downstream of the shock. This explains the difference between the two spectral index maps and in the right column of Fig. 3, which indicate the spectral slope of the radio flux density at low and high frequencies, respectively (). In these images we applied a radial mask, in order to reduce the noise in the outskirts (indicated by the dashed circle in the upper-left panel). The spectral slope at low frequencies is generally flatter (with maximum values of about at the shock front), than at high frequencies (about at shock), where the bulk of CR electrons has already cooled off to lower energies, causing a steepening of the spectrum. This aspect, as well as a comparison to observed counterparts, is described in Sect. 3.4.
@@ 3.4 Emission spectrum and total radio power
According to the formation model analysed in this paper, the radio feature that we identify as an ORC is a merger shock around a massive galaxy or galaxy group, triggered along the axis of a major merger. In this scenario, the total radio luminosity would be the highest shortly after the merger event, since that would be the time when the shock can inject relatively many CR electron populations efficiently into the dense medium at small radii. If we assume that CRs get accelerated from the thermal pool according to DSA, the power-law index of the injected CR energy spectrum depends only on the Mach number . Rewriting in terms of the spectral index of the emission spectrum yields
@@
In the limit of a strong shock (), the spectral index is , while weaker shocks generate steeper spectra (). After the initial central merger, the shock travels radially outwards with decreasing effective Mach numbers, producing increasingly steep spectra at the front.
@@
In Fig. 4 we present the time evolution of the ORC’s total luminosity produced by synchrotron emission. For this purpose, we extracted the particles around the expanding shock front at each time step and calculated their cumulative radio brightness at 50 frequencies equidistant in space according to Eq. A1 by Böss et al. (2023a). The two panels show the total brightness on the y-axis, while the x-axis represents time and frequency in the left and right panel, respectively. On the left, different colours identify varying frequencies according to the colour bar, while the solid black line highlights specifically , namely the frequency at which ORC Physalis (Koribalski et al. 2024b) was observed. The vertical dashed lines indicate the three times at which the simulation was displayed in Fig. 1 (). Shortly after , the merger triggers the shock – evident by the sudden jump in luminosity – and reaches its maximum by around . Afterwards, the total power steadily decreases in all frequencies, with the decline being stronger at higher frequencies. This effect becomes even more apparent on the right side of Fig. 4. This panel shows the cumulative emission spectrum of the simulated ORC at four successive times, which are approximately apart, starting from the time of maximum radio brightness at (yellow line). Early on, we see a relatively flat spectral index of in the lower frequency range. Moving on to later times, the spectrum rapidly decreases in total luminosity, with further spectral steepening at higher frequencies. This is due to two effects: while the shock travels outwards, its effective Mach number decreases, resulting in a lower acceleration efficiency in the DSA framework implemented in our code. Newly injected CRs therefore display increasingly steep spectra with time (see Eq. 1), while simultaneously the normalisation of the spectrum systematically decreases. The latter feature is also influenced by the decline in local gas density and magnetic field strength with radius, since these circumstances lead to less gas with new CR populations and lower synchrotron emissivity, respectively. For a fixed time step, the steeper slope at high frequencies is caused by radiation losses, since the energy loss timescale of electrons due to Synchrotron emission and inverse Compton scattering strongly depends on their energy. This leads to more rapid energy loss of CR electrons at higher energy bins, which is reflected in the steeper emission spectrum at higher frequencies, since the contribution to the spectral emission of a charged particle in a magnetic field is closely peaked around an energy-dependent maximum (Ginzburg 1979).
@@
Taking the spectral index at the earliest time () as a proxy for the overall maximum over time, we can easily estimate the respective Mach number by rearranging Eq. 1 for . This would yield a radio-inferred Mach number of at the ORC, which is comparable to measurements of merger events in galaxy clusters. The actual maximum of the Mach number distribution is lower (around ). However, the distribution is relatively broad (see Fig. 8 in the Appendix). The maximum radio power at is , which is about three orders of magnitude lower than the measurement for ORC Physalis (Koribalski et al. 2024b).
3.5 Polarisation
@@
Synchrotron radiation from electrons is polarised with a polarisation angle , i.e. the angle between the electric field vector of the associated radiation, and a chosen unit vector in the sky plane, defined as orthogonal to the line of sight. Since the polarisation angle is only dependent on local properties, the so-called polarisation fraction is relatively high in case of a singular, relatively small emitting CR cloud where the local magnetic field is approximately uniform. But as the number of contributing electrons along the line of sight increases, the fractional polarisation of the observed radiation consequently decreases because the direction of the magnetic field is not homogeneous. Therefore, a polarisation analysis of radio observations can be a powerful tool to dissect the physical processes relevant in ORCs. Since our simulation self-consistently evolves CR electrons and the magnetic field direction, we can provide polarisation information of the simulated ORC and compare it with observed counterparts for the first time.
@@
The polarisation state of electromagnetic radiation at a given frequency is commonly described by the four Stokes parameters and , since they can be inferred directly from observed quantities. Generally, synchrotron radiation is elliptically polarised. But in the limit of ultra-relativistic electrons, the light cone of radiation – defined as the solid angle around the velocity vector within which an observer receives a significant signal – becomes much smaller than . In this case, the observer receives only ’flashes’ of synchrotron radiation when the electron moves almost exactly towards the observer. The associated electric field vector hence only draws a line in the sky plane (instead of an ellipse), which is equivalent to linear polarisation (Ginzburg 1979; Longair 2011). In this limit, is vanishes and only three Stokes parameters remain. Assuming the line of sight is along the -direction, is defined as the total intensity received by the observer integrated along the line of sight:
@@
where is the synchrotron emissivity at a given frequency and position . Meanwhile and , which carry the polarisation information, are defined by
@@
where is the polarised synchrotron emissivity. Note that the polarisation angle is subject to Faraday rotation along the line of sight. The polarisation angle therefore needs to be adjusted at each contributing particle along the line of sight according to the travelled distance by
@@
where is the observed wavelength, RM the rotation measure, the local electron number density and the component of the magnetic field vector along the line of sight. The polarisation angle depends only on the magnetic field components in the sky plane and is given by
Finally, the polarisation fraction can be obtained via
@@
Generally, two factors will influence the degree of polarisation of synchrotron radiation: i) the energy spectrum of the underlying CR electron population and ii) the turbulence of the magnetic field and hence of the carrying intergalactic medium. In the case of a uniform magnetic field and a power law as the CR energy distribution, the polarisation fraction only depends on the power-law index of this energy distribution, which can be rewritten in terms of the spectral index of the emission spectrum (cf. Eq. 5.46 by Ginzburg 1979):
@@
It is instructive to keep this limit in mind, since it provides us with a sense of what to expect for a given CR population. If the measured value for the fractional polarisation is significantly lower, it is likely due to turbulence in the magnetic field along the line of sight. This produces an increasingly diverse distribution of polarisation angles, which can further ’fan out’ due to Faraday rotation.
In Fig. 5 we present the polarisation fraction of the simulated ORC at a frequency of given by Eq. 7. Faraday rotation along the line of sight is taken into account according to Eq. 5. Each quantity ( and ) was smoothed according to the resolution of ASKAP at (, see Sect. 3.3), where we assumed the same distance as to ORC Physalis at (Koribalski et al. 2024b). The time, scale, orientation and circular mask of the map is the same as in Fig. 3. For the most part, the simulated ORC has a degree of polarisation between . We analysed the rotation measure for the simulation and found that it is negligible outside the halo centre (see Fig. 9). Also, the beam size of 2.5 kpc adopted for the mock image is much smaller than the smoothing length of 10 kpc for the gas particles at the ORC position (see Sect. 3.6). This means that the polarisation is dominated by emission at the shock ring. We note that magnetic field fluctuations below our resolution limit must exist and might contribute to the final polarisation. However, is is not possible to make a quantitative prediction for this particular effect based on the this simulation.
@@
Judging by their emission spectrum, CRs are injected quite efficiently in our simulation and yield a spectral index close to the DSA limit of (see Sect. 3.4). The fact that we arrive at a much lower mean polarisation degree than the expected value given by Eq. 8 is an indicator for the turbulence that is present inside the shocked plasma along the line of sight, as explained above. The sporadic white spots in Fig. 5, which show a deviantly high degree of polarisation in comparison, correspond to sight lines with singular CR populations that are not contaminated by additional emissions along the line of sight. This is mostly a numerical artefact, which is also evident from the overall ’popcorn-like’ appearance of the map. Although we are already working with very high resolution here, the shock finder (which is highly resolution dependent) occasionally fails to pinpoint the shock location smoothly (see Sect. 3.6). Additionally, the injection of CRs strongly depends on the evaluated Mach number as well as on the estimated shock obliquity, which determines the local acceleration efficiency and can further add uncertainties to the simulated polarisation. We point out, however, that this has no significant impact on the overall value of the polarisation fraction across the shock ring. Fig. 5 therefore qualifies for a comparison with observed ORCs, which is carried out in the discussion (Sect. 4).